# Measurement of medium–induced acoplanarity in central $\mathrm{Au}\text{--}\mathrm{Au}$ and $pp$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV using direct–photon+jet and $\pi^{0}$+jet correlations

### v1

@@ Abstract
Introduction - Matter under extreme conditions of temperature and density forms a state of matter called Quark–Gluon Plasma (QGP) which consists of deconfined quarks and gluons [1, 2, 3, 4]. A QGP filled the early universe a few microseconds after the Big Bang, and is generated in collisions of heavy nuclei at the Relativistic Heavy–Ion Collider (RHIC) and the Large Hadron Collider (LHC). Experimental measurements at these facilities, and their comparison to theoretical model calculations, show that the QGP exhibits emergent collective behavior, flowing with the lowest possible specific shear viscosity [5].
@@
The secondary scattering of hard partons in the QGP has been proposed as a probe of QGP quasiparticles [19, 20, 21, 22, 23, 24], in analogy to Rutherford scattering as a probe of the atomic nucleus [25]. Observation of secondary, in–medium partonic scattering requires a coincidence observable, in which a trigger particle associated with an initial hard scattering specifies a direction, and the azimuthal difference $\Delta\phi$ between a recoil jet and the trigger (acoplanarity with respect to the plane defined by the beam axis and trigger) is measured. Acoplanarity distributions have been measured at the LHC for $pp$ and central $\mathrm{Pb}\text{--}\mathrm{Pb}$ collisions at center of mass energy $\sqrt{s_{\mathrm{NN}}}=5.02$ TeV, by the CMS and ALICE collaborations [26, 27, 28, 29, 30, 31]. CMS utilizes a direct–photon ($\gamma_{\mathrm{dir}}$) trigger with transverse energy $E_{\mathrm{T}}>60$ GeV and recoil jets with transverse momentum $p_{\mathrm{T,jet}}>30$ $\mathrm{GeV/}c$, with jet resolution parameter $R=0.3$; no medium–induced modification is observed within uncertainties [30]. ALICE reports semi-inclusive distributions of charged–particle recoil jets with $p_{\mathrm{T,jet}}^{\mathrm{ch}}>10$ $\mathrm{GeV/}c$ recoiling from hadron triggers with $p_{\mathrm{T}}^{\mathrm{trig}}>20$ $\mathrm{GeV/}c$ (h+jet); significant medium–induced broadening of the acoplanarity distribution is observed in the range $2010<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 20 $\mathrm{GeV/}c$ for $R=0.4$ and 0.5, but not for $R=0.2$ or at higher $p_{\mathrm{T,jet}}^{\mathrm{ch}}$ [27, 28]. This marked dependence of the broadening on $R$ and $p_{\mathrm{T,jet}}^{\mathrm{ch}}$ suggests that it arises from response of the QGP medium to excitation by a jet (“wake”), rather than single hard Rutherford–like scattering [27, 28]. At RHIC, the STAR collaboration has also reported an acoplanarity measurement based on semi–inclusive h+jet correlations in central and peripheral $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV, with no significant medium–induced broadening observed for $R=0.3$ [32].
@@
In this Letter, the STAR experiment reports the first measurement of jet acoplanarity of charged–particle jets recoiling from direct photon ($\gamma_{\mathrm{dir}}$) and $\pi^{0}$ triggers in $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV. The trigger particles have $1511<E_{\mathrm{T}}^{\mathrm{trig}}<$ 15 GeV, with the semi–inclusive distribution of recoil jets reported in $2010<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 20 $\mathrm{GeV/}c$ for $R=0.2$ and 0.5. Uncorrelated jet background yield is corrected using event mixing. This letter extends the analysis reported in Refs. [33, 34], to measure acoplanarity.
@@
Detector, dataset, and analysis - Data for $pp$ and $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV were recorded during the 2009 and 2014 RHIC runs, respectively. The detector, datasets, triggering, offline event selection, and track reconstruction are described in Ref. [33]. Online event selection is based on high–energy single showers measured in the Barrel Electromagnetic Calorimeter (BEMC) [37]. Centrality of $\mathrm{Au}\text{--}\mathrm{Au}$ collisions is determined offline using the uncorrected charged–particle multiplicity within pseudo–rapidity $|\eta|<0.5$; the 15% highest–multiplicity (“central”) $\mathrm{Au}\text{--}\mathrm{Au}$ collisions are selected for analysis. Events are further selected in both the $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ datasets by requiring the presence of a $\gamma_{\mathrm{dir}}$ or $\pi^{0}$ candidate with $1511<E_{\mathrm{T}}^{\mathrm{trig}}<$ 15 GeV. The integrated luminosity for the analysis is 23 pb-1 and 3.9 nb-1 for $pp$ and $\mathrm{Au}\text{--}\mathrm{Au}$ collisions, respectively.
High-$p_{\mathrm{T}}$ photon production in RHIC collisions arises from several sources [38, 39]: direct ($2\rightarrow 2$) production (Compton, annihilation), fragmentation, and hadronic decays. Discrimination of $\gamma$ and $\pi^{0}$–induced showers utilizes EM shower shape measured in the BEMC and its Shower Maximum detector (BSMD) [40, 33]. The purity of the resulting $\pi^{0}$–tagged population is estimated from simulation to be greater than 95% [34], while the photon–tagged population contains an admixture of $\pi^{0}$ and is labeled “$\gamma$–rich” ($\gamma_{\mathrm{rich}}$). The $\gamma_{\mathrm{dir}}$ fraction of the $\gamma_{\mathrm{rich}}$ population, which is determined using the measured rate of nearby correlated charged hadrons, depends on collision system and $E_{\mathrm{T}}^{\mathrm{trig}}$ and is in the range $\approx 40-80\%$ [34]. Correction based on this rate accounts for the hadronic decay component, and much but not all of the fragmentation photon contribution [40, 34]. This population is labeled $\gamma_{\mathrm{dir}}$.
@@
The measured recoil–jet yield in central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions has multiple contributions: correlated recoil jets from the same hard (high–${Q}^{2}$) scattering process which generates the trigger, corresponding to $Y(p_{\mathrm{T,jet}}^{\mathrm{ch}},\Delta\phi)$; jets arising from other partonic scattering processes, which are uncorrelated with the trigger (multiple partonic interactions, or MPI); and combinatorial jets arising from the random combination of tracks generated by soft (low–${Q}^{2}$) processes. The jet yield in central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV due to MPI is negligible [43]. However, the combinatorial yield in such collisions can be significant relative to the correlated signal, especially for large $R$ at low $p_{\mathrm{T,jet}}^{\mathrm{ch}}$.
@@
Correction of the raw recoil-jet yield to measure the $Y(p_{\mathrm{T,jet}}^{\mathrm{ch}},\Delta\phi)$ distribution is first carried out on 1-D $Y(p_{\mathrm{T,jet}}^{\mathrm{reco,ch}})$ raw distributions binned in raw $\Delta\phi$, following the procedures described in Ref. [43]. The corrected 1-D distributions as a function of $p_{\mathrm{T,jet}}^{\mathrm{ch}}$ are then combined to form a corrected 2-D distribution in ($p_{\mathrm{T,jet}}^{\mathrm{ch}}$,$\Delta\phi$) using weights that account for $\Delta\phi$-smearing due to residual background. The correction steps are outlined below, with detail provided for the $\Delta\phi$–smearing correction.
@@
Figure 1: $Y_{\mathrm{SE}}$, $Y_{\mathrm{MEnorm}}$ and $\widetilde{Y}_{\Delta\phi}$ distributions (Eq. 3) as a function of $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ from $\gamma_{\mathrm{rich}}$ triggers with $1511<E_{\mathrm{T}}^{\mathrm{trig}}<$ 15 GeV in central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions, for $R=0.2$ (left panels) and 0.5 (right panels). $\widetilde{Y}$($p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$) datapoints with a negative central value not shown. Sub-panels for each $R$: $2.52.2<\Delta\phi<$ 2.5 (left), $3.02.8<\Delta\phi<$ 3.0 (right). Upper panels show SE and MEnorm distributions; lower panels show the ratio SE/MEnorm. Insets show the ratio in the normalization region.
@@
Correction for the combinatorial jet yield in central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions is carried out using mixed events (ME) [43, 33], which are constructed by combining single tracks from multiple real events (“same events,” or SE) in each of 540 distinct bins in multiplicity, $z_{\mathrm{vtx}}$ (primary vertex beamline position), event plane orientation, and run–averaged luminosity [33]. Figure 1 shows examples of the SE and ME (normalized, see below) distributions, and their ratio, for central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions with $\gamma_{\mathrm{rich}}$ triggers and recoil jets with $R=0.2$ and 0.5, in the ranges $2.52.2<\Delta\phi<$ 2.5 and $3.02.8<\Delta\phi<$ 3.0. A negative value of $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ arises if $p_{\mathrm{T,jet}}^{\mathrm{raw}}$ is less than $\rho A_{\mathrm{jet}}$; for large negative values of $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ this occurs predominantly due to uncorrelated background, resulting in the same $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$–dependent shape of the $Y_{\mathrm{SE}}$ and $Y_{\mathrm{ME}}$ distributions in that region [26, 43, 27, 28, 33, 34].
The ME distribution is normalized to the SE distribution in the negative $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ region (MEnorm), which corrects the effective acceptance difference of the SE and ME populations due to displacement of uncorrelated jet candidates at low and negative $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ by hard, correlated jet candidates [26, 43]. The SE/ME yield ratio (lower panel insets) is independent of $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ within a few percent over a negative–$p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ range in which the yield itself varies by several orders of magnitude; for $R=0.5$, the range in $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ over which the ratio is flat within statistical uncertainty is limited. The extracted normalization factors have values between 0.9 and unity [33, 34]. The systematic dependence of the normalization factor on the upper bound of the normalization region is negligible [44].
@@
The symbol $\widetilde{Y}_{\Delta\phi}$ denotes the distribution as a function of $p_{\mathrm{T}}$ in bins of $\Delta\phi$, while $\widetilde{Y}_{p_{\mathrm{T}}}$ (used below) denotes the distribution as a function of $\Delta\phi$ in bins of $p_{\mathrm{T}}$. This data–driven, statistical correction for uncorrelated yield enables recoil–jet measurements in central $A\text{--}A$ collisions over broad phase space, including low $p_{\mathrm{T,jet}}$ and large $R$ [33, 34, 27, 28]. No ME–based correction is applied for $pp$ collisions due to small background yield, i.e. $\widetilde{Y}$($p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$)=$Y_{\mathrm{SE}}$($p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$).
@@
Instrumental effects generate negligible $\Delta\phi$–smearing relative to the analysis binning. The only significant $\Delta\phi$–smearing is due to spatial variation of uncorrelated background in central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions, which can modify the jet centroid direction. Correction for $\Delta\phi$–smearing is implemented bin–wise by a weight matrix $w(\Delta\phi_{\mathrm{true}},\Delta\phi_{\mathrm{meas}})$ that scales the measured $\widetilde{Y}_{\Delta\phi}(p_{\mathrm{T,jet}}^{\mathrm{ch}})$ distributions. The weights are determined by embedding detector–level PYTHIA-generated events for $pp$ collisions into real central $\mathrm{Au}\text{--}\mathrm{Au}$ events, with systematic uncertainty determined by varying the jet fragmentation model to mimic jet quenching effects. The systematic uncertainty of this correction is negligible [44].
@@
Results - Figure 2 shows corrected $\widetilde{Y}_{p_{\mathrm{T}}}$($\Delta\phi$) distributions for $\gamma_{\mathrm{dir}}$ and $\pi^{0}$ triggers in $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions, for $R=0.2$ and 0.5. The distributions fall steeply away from $\Delta\phi=\pi$, with greater yield for $R=0.5$ than for $R=0.2$. Figure 1, lower panels, shows greater yield for $R=0.5$ than for $R=0.2$ for $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}>0$ at large angles relative to $\Delta\phi=\pi$; this effect is therefore not generated predominantly by corrections. A similar effect is reported in Ref. [27].
Figure 2: Corrected $\widetilde{Y}_{p_{\mathrm{T}}}$ distributions as a function of $\Delta\phi$ for $\gamma_{\mathrm{dir}}$ and $\pi^{0}$ triggers with $1511<E_{\mathrm{T}}^{\mathrm{trig}}<$ 15 GeV, in $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV for $R=0.2$ (left) and $R=0.5$ (right). Upper: $\gamma_{\mathrm{dir}}$ trigger, $1510<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 15 $\mathrm{GeV/}c$; middle: $\pi^{0}$ trigger, $1510<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 15 $\mathrm{GeV/}c$; lower: $\pi^{0}$ trigger, $2015<p_{\mathrm{T,jet}}^{\mathrm{ch}}<$ 20 $\mathrm{GeV/}c$. Data are plotted at the spectrum-weighted bin coordinate except for low–statistics points, which are shown as 95% CL upper limits. Theoretical calculations for $pp$ collisions are described in the text.

### v2

# STAR Collaboration

@@ Abstract
I Introduction
Matter under extreme conditions of temperature and density forms a state of matter called Quark–Gluon Plasma (QGP) which consists of deconfined quarks and gluons [32, 48, 22, 40]. A QGP filled the early universe a few microseconds after the Big Bang, and is generated in collisions of heavy nuclei at the Relativistic Heavy–Ion Collider (RHIC) and the Large Hadron Collider (LHC). Experimental measurements at these facilities, and their comparison to theoretical model calculations, show that the QGP exhibits emergent collective behavior, flowing with the lowest possible specific shear viscosity [41].
@@ I Introduction
The secondary scattering of hard partons in the QGP has been proposed as a probe of QGP quasiparticles [15, 20, 35, 34, 31, 36], in analogy to Rutherford scattering as a probe of the atomic nucleus [46]. Observation of secondary, in–medium partonic scattering may be observable using coincidence channels, in which a trigger particle associated with an initial hard scattering specifies a direction, and the azimuthal difference $\Delta\phi$ between a recoil jet and the trigger (acoplanarity with respect to the plane defined by the beam axis and trigger) is measured. Acoplanarity distributions have been measured at the LHC for $pp$ and central $\mathrm{Pb}\text{--}\mathrm{Pb}$ collisions [10, 7, 6, 29, 30, 49, 50]. Measurements by the CMS Collaboration utilize a direct–photon ($\gamma_{\mathrm{dir}}$) trigger with transverse energy $E_{\mathrm{T}}>40$ GeV and recoil jets with transverse momentum $p_{\mathrm{T,jet}}>30$ $\mathrm{GeV/}c$, with jet resolution parameter $R=0.3$; no medium–induced modification is observed within uncertainties [30, 50]. Likewise, no acoplanarity broadening is observed by CMS in high-$p_{\mathrm{T}}$ dijet [28] and Z+jet [49] correlations.
The ALICE Collaboration reports semi-inclusive distributions of charged–particle recoil jets with $p_{\mathrm{T,jet}}^{\mathrm{ch}}>10$ $\mathrm{GeV/}c$ recoiling from hadron triggers with $p_{\mathrm{T}}^{\mathrm{trig}}>20$ $\mathrm{GeV/}c$ (h+jet); significant medium–induced broadening of the acoplanarity distribution is observed in the range $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<20$ $\mathrm{GeV/}c$ for $R=0.4$ and 0.5, but not for $R=0.2$ or at higher $p_{\mathrm{T,jet}}^{\mathrm{ch}}$ [7, 6]. This marked dependence of the broadening on $R$ and $p_{\mathrm{T,jet}}^{\mathrm{ch}}$ for such low-$p_{\mathrm{T}}$ jets suggests that it arises from response of the QGP medium to excitation by a jet (“wake”), rather than single hard Rutherford–like scattering [7, 6]. CMS also reports a medium-induced statistical excess in soft–particle yield at large angles relative to the axis of sub-leading jets in unbalanced di-jet pairs, though without their incorporation in jet reconstruction or measurement of acoplanarity [42]. At RHIC, the STAR collaboration has also reported an acoplanarity measurement based on semi–inclusive h+jet correlations in central and peripheral $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV, with no significant medium–induced broadening observed for $R=0.3$ [13]. A search for in–medium partonic scattering has also been carried out using jet substructure [8].
@@
In this article, the STAR experiment reports the first measurement of jet acoplanarity of charged–particle jets recoiling from direct photon ($\gamma_{\mathrm{dir}}$) and $\pi^{0}$ triggers in $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV. The trigger particles have $11<E_{\mathrm{T}}^{\mathrm{trig}}<15$ GeV, with the semi–inclusive distribution of recoil jets reported in $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<20$ $\mathrm{GeV/}c$ for $R=0.2$ and 0.5. Uncorrelated jet background yield is corrected using event mixing. This article extends the analysis reported in Refs. [5, 4], to measure acoplanarity.
@@
II Detector, dataset, and analysis
Data for $pp$ and $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV were recorded during the 2009 and 2014 RHIC runs, respectively. The detector, datasets, triggering, offline event selection, and track reconstruction are described in Ref. [5]. Online event selection is based on high–energy single showers measured in the Barrel Electromagnetic Calorimeter (BEMC) [18]. Centrality of $\mathrm{Au}\text{--}\mathrm{Au}$ collisions is determined offline using the uncorrected charged–particle multiplicity within pseudo–rapidity $|\eta|<0.5$; the 15% highest–multiplicity (“central”) $\mathrm{Au}\text{--}\mathrm{Au}$ collisions are selected for analysis. Events are further selected in both the $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ datasets by requiring the presence of a $\gamma_{\mathrm{dir}}$ or $\pi^{0}$ candidate with $11<E_{\mathrm{T}}^{\mathrm{trig}}<15$ GeV. The integrated luminosity for the analysis is 23 pb-1 and 3.9 nb-1 for $pp$ and $\mathrm{Au}\text{--}\mathrm{Au}$ collisions, respectively.
High-$p_{\mathrm{T}}$ photon production in RHIC collisions arises from several sources [45, 37]: direct ($2\rightarrow 2$) production (Compton, annihilation), fragmentation, and hadronic decays. Discrimination of $\gamma$ and $\pi^{0}$–induced showers utilizes EM shower shape measured in the BEMC and its Shower Maximum detector (BSMD) [11, 5]. The purity of the resulting $\pi^{0}$–tagged population is estimated from simulation to be greater than 95% [4], while the photon–tagged population contains an admixture of $\pi^{0}$ and is labeled $\gamma$–rich” ($\gamma_{\mathrm{rich}}$). The $\gamma_{\mathrm{dir}}$ fraction of the $\gamma_{\mathrm{rich}}$ population, which is determined using the measured rate of nearby correlated charged hadrons, depends on collision system and $E_{\mathrm{T}}^{\mathrm{trig}}$ and is in the range $\approx 40-80\%$ [4]. Correction based on this rate accounts for the hadronic decay component, and much but not all of the fragmentation photon contribution [11, 4]. This population is labeled $\gamma_{\mathrm{dir}}$.
@@ II Detector, dataset, and analysis
[…]
@@
Correction of the raw recoil-jet yield to measure the $Y(p_{\mathrm{T,jet}}^{\mathrm{ch}},\Delta\phi)$ distribution is first carried out on 1-D $Y(p_{\mathrm{T,jet}}^{\mathrm{reco,ch}})$ raw distributions binned in raw $\Delta\phi$, following the procedures described in Ref. [13]. The corrected 1-D distributions as a function of $p_{\mathrm{T,jet}}^{\mathrm{ch}}$ are then combined to form a corrected 2-D distribution in ($p_{\mathrm{T,jet}}^{\mathrm{ch}}$$\Delta\phi$) using weights that account for $\Delta\phi$-smearing due to residual background. The correction steps are outlined below, with detail provided for the $\Delta\phi$–smearing correction.
@@
Figure 1: $Y_{\mathrm{SE}}$, $Y_{\mathrm{MEnorm}}$ and $\widetilde{Y}_{\Delta\phi}$ distributions (Eq. 3) as a function of $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ from $\gamma_{\mathrm{rich}}$ triggers with $11<E_{\mathrm{T}}^{\mathrm{trig}}<15$ GeV in central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions, for $R=0.2$ (top panels) and 0.5 (bottom panels). $\widetilde{Y}$($p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$) datapoints with a negative central value not shown. Sub-panels for each $R$: $2.2<\Delta\phi<2.5$ (left), $2.8<\Delta\phi<3.0$ (right). Upper panels show SE and MEnorm distributions; lower panels show the ratio SE/MEnorm. Insets show the ratio in the normalization region.
@@
Correction for the combinatorial jet yield in central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions is carried out using mixed events (ME) [13, 5], which are constructed by combining single tracks from multiple real events (“same events,” or SE) in each of 540 distinct bins in multiplicity, $z_{\mathrm{vtx}}$ (primary vertex beamline position), event plane orientation, and run–averaged luminosity [5]. Figure 1 shows examples of the SE and ME (normalized, see below) distributions, and their ratio, for central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions with $\gamma_{\mathrm{rich}}$ triggers and recoil jets with $R=0.2$ and 0.5, in the ranges $2.2<\Delta\phi<2.5$ and $2.8<\Delta\phi<3.0$. A negative value of $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ arises if $p_{\mathrm{T,jet}}^{\mathrm{raw}}$ is less than $\rho A_{\mathrm{jet}}$; for large negative values of $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$ this occurs predominantly due to uncorrelated background, resulting in the same $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}$–dependent shape of the $Y_{\mathrm{SE}}$ and $Y_{\mathrm{ME}}$ distributions in that region [10, 13, 7, 6, 5, 4].
[…]
@@
[…]
@@
Instrumental effects generate negligible $\Delta\phi$–smearing relative to the analysis binning. The only significant $\Delta\phi$–smearing is due to spatial variation of uncorrelated background in central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions, which can modify the jet centroid direction. Correction for $\Delta\phi$–smearing is implemented bin–wise by a weight matrix $w(\Delta\phi_{\mathrm{true}},\Delta\phi_{\mathrm{meas}})$ that scales the measured $\widetilde{Y}_{\Delta\phi}(p_{\mathrm{T,jet}}^{\mathrm{ch}})$ distributions. The weights are determined by embedding detector–level PYTHIA-generated events for $pp$ collisions into real central $\mathrm{Au}\text{--}\mathrm{Au}$ events, with systematic uncertainty determined by varying the jet fragmentation model to mimic jet quenching effects. The systematic uncertainty of this correction is negligible as discussed in the appendix.
@@
III Results
Figure 2 shows corrected $\widetilde{Y}_{p_{\mathrm{T}}}$($\Delta\phi$) distributions for $\gamma_{\mathrm{dir}}$ and $\pi^{0}$ triggers in $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions, for $R=0.2$ and 0.5. The distributions fall steeply away from $\Delta\phi=\pi$, with greater yield for $R=0.5$ than for $R=0.2$. Figure 1, lower panels, shows greater yield for $R=0.5$ than for $R=0.2$ for $p_{\mathrm{T,jet}}^{\mathrm{reco,ch}}>0$ at large angles relative to $\Delta\phi=\pi$; this effect is therefore not generated predominantly by corrections. A similar effect is reported in Ref. [7].
Figure 2: Corrected $\widetilde{Y}_{p_{\mathrm{T}}}$ distributions as a function of $\Delta\phi$ for $\gamma_{\mathrm{dir}}$ and $\pi^{0}$ triggers with $11<E_{\mathrm{T}}^{\mathrm{trig}}<15$ GeV, in $pp$ and central $\mathrm{Au}\text{--}\mathrm{Au}$ collisions at $\sqrt{s_{\mathrm{NN}}}=200$ GeV for $R=0.2$ (left) and $R=0.5$ (right). Upper: $\gamma_{\mathrm{dir}}$ trigger, $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$; middle: $\pi^{0}$ trigger, $10<p_{\mathrm{T,jet}}^{\mathrm{ch}}<15$ $\mathrm{GeV/}c$; lower: $\pi^{0}$ trigger, $15<p_{\mathrm{T,jet}}^{\mathrm{ch}}<20$ $\mathrm{GeV/}c$. Data are plotted at the spectrum-weighted bin coordinate except for low–statistics points, which are shown as 95% CL upper limits. Theoretical calculations for $pp$ collisions are described in the text.